Dr. Glendon, the school physician in charge of giving sports phyisicals, has compiled his information and has determined that the probability a student will be on a sports team is 0.39. yesterday. Dr. glendon examined five students chosen at random.
A)find to the nearest hundredth, the probability that at least 4 of the 5 students will be on a team.
B) find to the nearest hundredth, the probability that exactly one of the five students will not be on a team.

Answer by stanbon(75887)

Dr. Glendon, the school physician in charge of giving sports phyisicals, has compiled his information and has determined that the probability a student will be on a sports team is 0.39. yesterday. Dr. glendon examined five students chosen at random.
P(willl be on) = 0.39 ; P(will not be on) = 0.61
------------------------------
A)find to the nearest hundredth, the probability that at least 4 of the 5 students will be on a team.
P(4<= x <=5) = 1-P(0<= x <=3) = 1-binomcdf(5,0.39,3) = 0.0796
--------------------------------------
B) find to the nearest hundredth, the probability that exactly one of the five students will not be on a team.
P(x = 1) = 5C1(0.61)*(0.39)^4 = binompdf(5,0.61,1) = 0.0706
================
Cheers,
Stan H.